# Diffusion Model-Based Image Editing: A Survey

### v1

@@ 1 Introduction
In the realm of AI-generated content (AIGC), where artificial intelligence is leveraged to create and modify digital content, image editing is recognized as a significant area of innovation and practical applications. Different from image generation that creates new images from minimal inputs, image editing involves altering the appearance, structure, or content of an image, encompassing a range of changes from subtle adjustments to major transformations. This research is fundamental in various areas, including digital media, advertising, and scientific research, where altering visual content is essential. The evolution of image editing has reflected the advancements in digital technology, progressing from manual, labor-intensive processes to advanced digital techniques driven by learning-based algorithms. A pivotal advancement in this evolution is the introduction of Generative Adversarial Networks (GANs) [1, 2, 3, 4, 5, 6], which significantly enhance the possibilities for creative image manipulation.
More recently, diffusion models have emerged in AIGC [7, 8, 9, 10, 11, 12, 13, 14, 1, 15], leading to notable breakthroughs in visual generation tasks. Diffusion models, inspired by principles from non-equilibrium thermodynamics [15], work by gradually adding noise to data and then learning to reverse this process from random noise until generating desired data matching the source data distribution. They can be roughly classified into denoising diffusion based [15, 16, 17, 18] and score-matching based [19, 20, 21, 22, 23]. Their adaptability and effectiveness have led to widespread applications across various tasks such as image generation [24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38], video generation [39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56], image restoration [57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71], and image editing.
The application of diffusion models in image editing has experienced a surge in interest, evidenced by the significant number of research publications in this field over recent years. This growing attention highlights the potential and versatility of diffusion models in improving image editing performance compared to previous works. Given this significant advancement, it is essential to systematically review and summarize these contributions. However, existing survey literature on diffusion models concentrates on other specific visual tasks [72, 73, 74, 75], such as video applications [73] or image restoration and enhancement [74, 75]. Some surveys that mention image editing tend to provide only a cursory overview [76, 77, 78, 79, 80, 81, 82, 83], missing a detailed and focused exploration of the methods.
@@
To bridge this gap, we conduct a survey to offer an in-depth and comprehensive analysis that is distinctly focused on image editing. We delve deeply into the methodologies, input conditions, and a wide range of editing tasks achieved by diffusion models in this field. The survey critically reviews over 100 research papers, organizing them into three primary classes in terms of learning strategies: training-based approaches, testing-time finetuning approaches, and training and finetuning free approaches. Each class is further divided based on its core techniques, with detailed discussions presented in Sections 45, and 6, respectively. We also explore 10 distinct types of input conditions used in these methods, including text, mask, reference (Ref.) image, class, layout, pose, sketch, segmentation (Seg.) map, audio, and dragging points to show the adaptability of diffusion models in diverse image editing scenarios. Furthermore, our survey presents a new classification of image editing tasks into three broad categories, semantic editing, stylistic editing, and structural editing, covering 12 specific types. Fig. 1 visually represents the statistical distributions of research across learning strategies, input conditions, and editing task categories. In addition, we pay special attention to inpainting and outpainting, which together stand out as a unique type of editing. We explore both earlier traditional and current multimodal conditional methods, offering a comprehensive analysis of their methodologies in Section 7. We also introduce EditEval, a benchmark designed to evaluate text-guided image editing algorithms, as detailed in Section 8. In particular, an effective evaluation metric, LMM Score, is proposed by leveraging the advanced visual-language understanding capabilities of large multimodal models (LMMs). Finally, we present some current challenges and potential future trends as an outlook in Section 9.
@@ 2.1 Diffusion Models
where the set of hyper-parameters $0<\beta_{1:T}<1$ is indicative of the noise variance introduced at each sequential step. This diffusion process can be briefly expressed via a single-step equation:
@@
[…]
@@ 2.2.1 Conditional Image Generation
While our primary focus is on diffusion models in image editing, it is important to acknowledge related areas like conditional image generation. Unlike image editing which involves altering certain parts of an existing image, conditional image generation is about creating new images from scratch, guided by specified conditions. Early efforts [32, 31, 87, 88, 89, 90] typically involve class-conditioned image generation, which usually incorporates the class-induced gradients via an additional pretrained classifier during sampling. However, Ho et al. [86] introduce the classifier-free guidance, which does not rely on an external classifier and allows for more versatile conditions, e.g., text, as guidance.
Text-to-Image (T2I) Generation. GLIDE [34] is the first work that uses text to guide image generation directly from the high-dimensional pixel level, replacing the label in class-conditioned diffusion models. Similarly, Imagen [27] uses a cascaded framework to generate high-resolution images more efficiently in pixel space. A different line of research first projects the image into a lower-dimensional space and then applies diffusion models in this latent space. Representative works include Stable Diffusion (SD) [26], VQ-diffusion [91], and DALL-E 2 [25]. Following these pioneering studies, a large number of works [37, 92, 93, 94, 95, 96, 97] are proposed, advancing this field over the past two years.
Additional Conditions. Beyond text, more specific conditions are also used to achieve higher fidelity and more precise control in image synthesis. GLIGEN [98] inserts a gated self-attention layer between the original self-attention and cross-attention layers in each block of a pretrained T2I diffusion model for generating images conditioned on grounding boxes. Make-A-Scene [99] and SpaText [100] use segmentation masks to guide the generation process. In addition to segmentation maps, ControlNet [101] can also incorporate other types of input, such as depth maps, normal maps, canny edges, pose, and sketches as conditions. Other methods like UniControlNet [102], UniControl [103], Composer [104], and T2I-Adapter [105] integrate diverse conditional inputs and include additional layers, enhancing the generative process controlled by these conditions.
@@
Customized Image Generation. Closely related to image editing within conditional image generation is the task of creating personalized images. This task focuses on generating images that maintain a certain identity, typically guided by a few reference images of the same subject. Two early approaches addressing this customized generation through few-shot images are Textual Inversion [106] and DreamBooth [107]. Specifically, Textual Inversion learns a unique identifier word to represent a new subject and incorporates this word into the text encoder’s dictionary. DreamBooth, on the other hand, binds a new, rare word with a specific subject by finetuning the entire Imagen [27] model with a few reference images. To combine multiple new concepts effectively, CustomDiffusion [108] only optimizes the cross-attention parameters in Stable Diffusion [26], representing new concepts and conducting joint training for the multiple-concept combination. Following these foundational works, subsequent methods [109, 110, 111, 112, 113, 114, 115, 116, 117] provide more refined control over the generated images, enhancing the precision and accuracy of the output.
@@ 2.2.2 Image Restoration and Enhancement
Method Condition(s) Semantic Editing Stylistic Editing Structural Editing
Obj. Add. Obj. Remo. Obj. Repl. Bg. Chg. Emo. Expr. Mod. Color Chg. Style Chg. Text. Chg. Obj. Move. Obj. Size. Chg. Obj. Act. Chg. Persp. /View. Chg.
DiffusionCLIP  [131] Text, Class
Asyrp  [132] Text, Class
EffDiff  [133] Class
DiffStyler  [134] Text, Class
StyleDiffusion  [135] Ref. Image, Class
UNIT-DDPM  [136] Class
CycleNet  [137] Class
Diffusion Autoencoders  [138] Class
HDAE  [139] Class
EGSDE  [140] Class
Pixel-Guided Diffusion  [141] Seg. Map, Class
PbE  [142] Ref. Image
RIC  [143] Ref. Image, Sketch
ObjectStitch  [144] Ref. Image
PhD  [145] Ref. Image, Layout
DreamInpainter  [146] Ref. Image, Mask, Text
Anydoor  [147] Ref. Image, Mask
FADING  [148] Text
PAIR Diffusion  [149] Ref. Image, Text
SmartBrush  [150] Text, Mask
IIR-Net  [151] Text
PowerPaint  [152] Text, Mask
Imagen Editor  [153] Text, Mask
SmartMask  [154] Text
Uni-paint  [155] Text, Mask, Ref. Image
InstructPix2Pix  [156] Text
MoEController  [157] Text
FoI  [158] Text
LOFIE  [159] Text
InstructDiffusion  [160] Text
Emu Edit  [161] Text
DialogPaint  [162] Text
Inst-Inpaint  [163] Text
HIVE  [164] Text
ImageBrush  [165] Ref. Image, Seg. Map, Pose
InstructAny2Pix  [166] Ref. Image, Audio, Text
MGIE  [167] Text
SmartEdit  [168] Text
iEdit  [169] Text
TDIELR  [170] Text
ChatFace  [171] Text
UniTune  [172] Text
Custom-Edit  [173] Text, Ref. Image
KV-Inversion  [174] Text
Null-Text Inversion  [175] Text
DPL  [176] Text
DiffusionDisentanglement  [177] Text
Prompt Tuning Inversion  [178] Text
@@ 3 Categorization of Image Editing
[…]
Training-Based
Approaches
Domain-Specific Editing with Weak Supervision DiffusionCLIP [131], Asyrp [132], EffDiff [133], DiffStyler [134], StyleDiffusion [135], UNIT-DDPM [136], CycleNet [137], Diffusion Autoencoders [138], HDAE [139], EGSDE [140], Pixel-Guided Diffusion [141] Reference and Attribute Guidance via Self-Supervision PbE [142], RIC [143], ObjectStitch [144], PhD [145], DreamInpainter [146], Anydoor [147], FADING [148], PAIR Diffusion [149], SmartBrush [150], IIR-Net [151], PowerPaint [152], Imagen Editor [153], SmartMask [154], Uni-paint [155] Instructional Editing via Full Supervision InstructPix2Pix [156], MoEController [157], FoI [158], LOFIE [159], InstructDiffusion [160], Emu Edit [161], DialogPaint [162], Inst-Inpaint [163], HIVE [164], ImageBrush [165], InstructAny2Pix [166], MGIE [167], SmartEdit [168] Pseudo-Target Retrieval with Weak Supervision iEdit [169], TDIELR [170], ChatFace [171]
@@ 4.1 Domain-Specific Editing with Weak Supervision
Figure 3: Comparison of the training pipelines between two representative CLIP guided methods: DiffusionCLIP [131] and Asyrp [132]. The sample images are from Asyrp [132] on the CelebA [236] dataset.
@@ 5.1 Denoising Model Finetuning
[…]
Testing-Time
Finetuning Approaches
Denosing Model Finetuning UniTune [172], Custom-Edit [173], KV-Inversion [174] Embedding Finetuning Null-Text Inversion [175], DPL [176], DiffusionDisentanglement [177], Prompt Tuning Inversion [178] Guidance with a Hypernetwork StyleDiffusion [179], InST [180] Latent Variable Optimization DragonDiffusion [181], DragDiffusion [182], DDS [183], DiffuseIT [184], CDS [185], MagicRemover [186] Hybrid Finetuning Imagic [187], LayerDiffusion [188], Forgedit [189], SINE [190]
@@ 5.5 Hybrid Finetuning
[…]
Training and Finetuning
Free Approaches
Input Text Refinement PRedItOR [191], ReDiffuser [192], Captioning and Injection [193], InstructEdit [194] Inversion/Sampling Modification Direct Inversion [195], DDPM Inversion [196], SDE-Drag [197], LEDITS++ [198], FEC [199], EMILIE [200], Negative Inversion [201], Proximal Inversion [202], Null-Text Guidance [203], EDICT [204], AIDI [205], CycleDiffusion [206], InjectFusion [207], Fixed-point inversion [208], TIC [209], Diffusion Brush [210], Self-guidance [211] Attention Modification P2P [212], Pix2Pix-Zero [213], MasaCtrl [85], PnP [214], TF-ICON [215], Object-Shape Variations [216], Conditional Score Guidance [217], EBMs [218], Shape-Guided Diffusion [219], HD-Painter [220] Mask Guidance FISEdit [221], Blended Latent Diffusion [222], PFB-Diff [223], DiffEdit [224], RDM [225], MFL [226], Differential Diffusion [227], Watch Your Steps [228], Blended Diffusion [229], ZONE [230], Inpaint Anything [231] Multi-Noise Redirection The Stable Artist [232], SEGA [233], LEDITS [234], OIR-Diffusion [235]

### v2

@@ 1 Introduction
In the realm of AI-generated content (AIGC), where artificial intelligence is leveraged to create and modify digital content, image editing is recognized as a significant area of innovation and practical applications. Different from image generation that creates new images from minimal inputs, image editing involves altering the appearance, structure, or content of an image, encompassing a range of changes from subtle adjustments to major transformations. This research is fundamental in various areas, including digital media, advertising, and scientific research, where altering visual content is essential. The evolution of image editing has reflected the advancements in digital technology, progressing from manual, labor-intensive processes to advanced digital techniques driven by learning-based algorithms. A pivotal advancement in this evolution is the introduction of Generative Adversarial Networks (GANs) [1, 2, 3, 4, 5, 6], which significantly enhance the possibilities for creative image manipulation. More recently, diffusion models have emerged in AIGC [7, 8, 9, 10, 11, 12, 13, 14, 1, 15], leading to notable breakthroughs in visual generation tasks. Diffusion models, inspired by principles from non-equilibrium thermodynamics [15], work by gradually adding noise to data and then learning to reverse this process from random noise until generating desired data matching the source data distribution. They can be roughly classified into denoising diffusion based [15, 16, 17, 18] and score-matching based [19, 20, 21, 22, 23]. Their adaptability and effectiveness have led to widespread applications across various tasks such as image generation [24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38], video generation [39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56], image restoration [57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71], and image editing.
@@
The application of diffusion models in image editing has experienced a surge in interest, evidenced by the significant number of research publications in this field over recent years. This growing attention highlights the potential and versatility of diffusion models in improving image editing performance compared to previous works. Given this significant advancement, it is essential to systematically review and summarize these contributions. However, existing survey literature on diffusion models concentrates on other specific visual tasks [72, 73, 74, 75], such as video applications [73] or image restoration and enhancement [74, 75]. Some surveys that mention image editing tend to provide only a cursory overview [76, 77, 78, 79, 80, 81, 82, 83], missing a detailed and focused exploration of the methods. To bridge this gap, we conduct a survey to offer an in-depth and comprehensive analysis that is distinctly focused on image editing. We delve deeply into the methodologies, input conditions, and a wide range of editing tasks achieved by diffusion models in this field. The survey critically reviews over 100 research papers, organizing them into three primary classes in terms of learning strategies: training-based approaches, testing-time finetuning approaches, and training and finetuning free approaches. Each class is further divided based on its core techniques, with detailed discussions presented in Sections 45, and 6, respectively. We also explore 10 distinct types of input conditions used in these methods, including text, mask, reference (Ref.) image, class, layout, pose, sketch, segmentation (Seg.) map, audio, and dragging points to show the adaptability of diffusion models in diverse image editing scenarios. Furthermore, our survey presents a new classification of image editing tasks into three broad categories, semantic editing, stylistic editing, and structural editing, covering 12 specific types. Fig. 1 visually represents the statistical distributions of research across learning strategies, input conditions, and editing task categories. In addition, we pay special attention to inpainting and outpainting, which together stand out as a unique type of editing. We explore both earlier traditional and current multimodal conditional methods, offering a comprehensive analysis of their methodologies in Section 7. We also introduce EditEval, a benchmark designed to evaluate text-guided image editing algorithms, as detailed in Section 8. In particular, an effective evaluation metric, LMM Score, is proposed by leveraging the advanced visual-language understanding capabilities of large multimodal models (LMMs). Finally, we present some current challenges and potential future trends as an outlook in Section 9.
@@ 2.1 Diffusion Models
where the set of hyper-parameters $10<\beta_{1:T}<$ 1 is indicative of the noise variance introduced at each sequential step. This diffusion process can be briefly expressed via a single-step equation:
@@
In the realm of text-conditional image generation, it is vital to ensure substantial textual influence and control over the generated output. To this end, Ho et al. [86] introduce the concept of classifier-free guidance, a technique that amalgamates conditional and unconditional predictions. More specifically, let $\varnothing=\psi(``")$ [^1] The placeholder $\varnothing$ is typically used for negative prompts to prevent certain attributes from manifesting in the generated image. denote the null text embedding. When combined with a guidance scale $w$, the classifier-free guidance prediction is formalized as:
@@ 2.2.1 Conditional Image Generation
While our primary focus is on diffusion models in image editing, it is important to acknowledge related areas like conditional image generation. Unlike image editing which involves altering certain parts of an existing image, conditional image generation is about creating new images from scratch, guided by specified conditions. Early efforts [32, 31, 87, 88, 89, 90] typically involve class-conditioned image generation, which usually incorporates the class-induced gradients via an additional pretrained classifier during sampling. However, Ho et al. [86] introduce the classifier-free guidance, which does not rely on an external classifier and allows for more versatile conditions, e.g., text, as guidance. Text-to-Image (T2I) Generation. GLIDE [34] is the first work that uses text to guide image generation directly from the high-dimensional pixel level, replacing the label in class-conditioned diffusion models. Similarly, Imagen [27] uses a cascaded framework to generate high-resolution images more efficiently in pixel space. A different line of research first projects the image into a lower-dimensional space and then applies diffusion models in this latent space. Representative works include Stable Diffusion (SD) [26], VQ-diffusion [91], and DALL-E 2 [25]. Following these pioneering studies, a large number of works [37, 92, 93, 94, 95, 96, 97] are proposed, advancing this field over the past two years. Additional Conditions. Beyond text, more specific conditions are also used to achieve higher fidelity and more precise control in image synthesis. GLIGEN [98] inserts a gated self-attention layer between the original self-attention and cross-attention layers in each block of a pretrained T2I diffusion model for generating images conditioned on grounding boxes. Make-A-Scene [99] and SpaText [100] use segmentation masks to guide the generation process. In addition to segmentation maps, ControlNet [101] can also incorporate other types of input, such as depth maps, normal maps, canny edges, pose, and sketches as conditions. Other methods like UniControlNet [102], UniControl [103], Composer [104], and T2I-Adapter [105] integrate diverse conditional inputs and include additional layers, enhancing the generative process controlled by these conditions. Customized Image Generation. Closely related to image editing within conditional image generation is the task of creating personalized images. This task focuses on generating images that maintain a certain identity, typically guided by a few reference images of the same subject. Two early approaches addressing this customized generation through few-shot images are Textual Inversion [106] and DreamBooth [107]. Specifically, Textual Inversion learns a unique identifier word to represent a new subject and incorporates this word into the text encoder’s dictionary. DreamBooth, on the other hand, binds a new, rare word with a specific subject by finetuning the entire Imagen [27] model with a few reference images. To combine multiple new concepts effectively, CustomDiffusion [108] only optimizes the cross-attention parameters in Stable Diffusion [26], representing new concepts and conducting joint training for the multiple-concept combination. Following these foundational works, subsequent methods [109, 110, 111, 112, 113, 114, 115, 116, 117] provide more refined control over the generated images, enhancing the precision and accuracy of the output.
@@ 2.2.2 Image Restoration and Enhancement
Method
Condition(s)
Semantic Editing Stylistic Editing Structural Editing
Obj. Add.
Obj. Remo.
Obj. Repl.
Bg. Chg.
Emo. Expr. Mod.
Color Chg.
Style Chg.
Text. Chg.
Obj. Move.
Obj. Size. Chg.
Obj. Act. Chg.
Persp. /View. Chg.
DiffusionCLIP  [131]
Text, Class
Asyrp  [132]
Text, Class
EffDiff  [133]
Class
DiffStyler  [134]
Text, Class
StyleDiffusion  [135]
Ref. Image, Class
UNIT-DDPM  [136]
Class
CycleNet  [137]
Class
Diffusion Autoencoders  [138]
Class
HDAE  [139]
Class
EGSDE  [140]
Class
Pixel-Guided Diffusion  [141]
Seg. Map, Class
PbE  [142]
Ref. Image
RIC  [143]
Ref. Image, Sketch
ObjectStitch  [144]
Ref. Image
PhD  [145]
Ref. Image, Layout
DreamInpainter  [146]
Ref. Image, Mask, Text
Anydoor  [147]
Ref. Image, Mask
FADING  [148]
Text
PAIR Diffusion  [149]
Ref. Image, Text
SmartBrush  [150]
Text, Mask
IIR-Net  [151]
Text
PowerPaint  [152]
Text, Mask
Imagen Editor  [153]
Text, Mask
SmartMask  [154]
Text
Uni-paint  [155]
Text, Mask, Ref. Image
InstructPix2Pix  [156]
Text
MoEController  [157]
Text
FoI  [158]
Text
LOFIE  [159]
Text
InstructDiffusion  [160]
Text
Emu Edit  [161]
Text
DialogPaint  [162]
Text
Inst-Inpaint  [163]
Text
HIVE  [164]
Text
ImageBrush  [165]
Ref. Image, Seg. Map, Pose
InstructAny2Pix  [166]
Ref. Image, Audio, Text
MGIE  [167]
Text
SmartEdit  [168]
Text
iEdit  [169]
Text
TDIELR  [170]
Text
ChatFace  [171]
Text
UniTune  [172]
Text
Custom-Edit  [173]
Text, Ref. Image
KV-Inversion  [174]
Text
Null-Text Inversion  [175]
Text
DPL  [176]
Text
DiffusionDisentanglement  [177]
Text
Prompt Tuning Inversion  [178]
Text
@@ 3 Categorization of Image Editing
{forest}
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hidden-draw, rounded corners, align=left, minimum width=2em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=4.5em,font=,, where level=2text width=3.1em,font=,, where level=3text width=3.0em,font=,, [ Training-Based
Approaches [ Domain-Specific Editing
with Weak Supervision, text width=7.5em [ DiffusionCLIP [131], Asyrp [132], EffDiff [133], DiffStyler [134], StyleDiffusion [135],
UNIT-DDPM [136], CycleNet [137], Diffusion Autoencoders [138], HDAE [139],
EGSDE [140], Pixel-Guided Diffusion [141] , leaf, text width=27em ] ] [ Reference and Attribute
Guidance via
Self-Supervision, text width=7.5em [ PbE [142], RIC [143], ObjectStitch [144], PhD [145], DreamInpainter [146],
Anydoor [147], FADING [148], PAIR Diffusion [149], SmartBrush [150],
IIR-Net [151], PowerPaint [152], Imagen Editor [153], SmartMask [154],
Uni-paint [155] , leaf, text width=24em ] ] [ Instructional Editing
via Full Supervision, text width=6.5em [ InstructPix2Pix [156], MoEController [157], FoI [158], LOFIE [159],
InstructDiffusion [160], Emu Edit [161], DialogPaint [162], Inst-Inpaint [163],
HIVE [164], ImageBrush [165], InstructAny2Pix [166], MGIE [167], SmartEdit [168] , leaf, text width=27em ] ] [ Pseudo-Target Retrieval
with Weak Supervision, text width=7.5em [ iEdit [169], TDIELR [170], ChatFace [171] , leaf, text width=14em ] ] ]
@@ 4.1 Domain-Specific Editing with Weak Supervision
Figure 3: Comparison of the training pipelines between two representative CLIP guided methods: DiffusionCLIP [131] and Asyrp [132]. The sample images are from DiffusionCLIP [131] and Asyrp [132] on the AFHQ [237] dataset.
@@ 5.1 Denoising Model Finetuning
{forest}
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hidden-draw, rounded corners, align=left, minimum width=2em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=4.5em,font=,, where level=2text width=3.1em,font=,, where level=3text width=3.0em,font=,, [ Testing-Time
Finetuning Approaches [ Denosing Model
Finetuning, text width=5.5em [ UniTune [172], Custom-Edit [173], KV-Inversion [174] , leaf, text width=17em ] ] [ Embedding
Finetuning, text width=4em [ Null-Text Inversion [175], DPL [176], DiffusionDisentanglement [177],
Prompt Tuning Inversion [178] , leaf, text width=22.4em ] ] [ Guidance with
a Hypernetwork, text width=5em [ StyleDiffusion [179], InST [180] , leaf, text width=10.5em ] ] [ Latent Variable
Optimization, text width=5em [ DragonDiffusion [181], DragDiffusion [182], DDS [183],
DiffuseIT [184], CDS [185], MagicRemover [186] , leaf, text width=18em ] ] [ Hybrid
Finetuning, text width=3.5em [ Imagic [187], LayerDiffusion [188], Forgedit [189], SINE [190] , leaf, text width=20em ] ] ]
@@ 5.5 Hybrid Finetuning
{forest}
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hidden-draw, rounded corners, align=left, minimum width=2em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=4.5em,font=,, where level=2text width=3.1em,font=,, where level=3text width=3.0em,font=,, [ Training and Finetuning
Free Approaches [ Input Text
Refinement, text width=4em [ PRedItOR [191], ReDiffuser [192], Captioning and Injection [193], InstructEdit [194] , leaf, text width=26em ] ] [ Inversion/Sampling
Modification, text width=6em [ Direct Inversion [195], DDPM Inversion [196], SDE-Drag [197], LEDITS++ [198],
FEC [199], EMILIE [200], Negative Inversion [201], Proximal Inversion [202],
Null-Text Guidance [203], EDICT [204], AIDI [205], CycleDiffusion [206],
InjectFusion [207], Fixed-point inversion [208], TIC [209], Diffusion Brush [210],
Self-guidance [211] , leaf, text width=25.5em ] ] [ Attention
Modification, text width=4em [ P2P [212], Pix2Pix-Zero [213], MasaCtrl [85], PnP [214], TF-ICON [215],
Object-Shape Variations [216], Conditional Score Guidance [217], EBMs [218],
Shape-Guided Diffusion [219], HD-Painter [220] , leaf, text width=25em ] ] [ Mask
Guidance, text width=3em [ FISEdit [221], Blended Latent Diffusion [222], PFB-Diff [223], DiffEdit [224], RDM [225],
MFL [226], Differential Diffusion [227], Watch Your Steps [228], Blended Diffusion [229],
ZONE [230], Inpaint Anything [231] , leaf, text width=28em ] ] [ Multi-Noise
Redirection, text width=4em [ The Stable Artist [232], SEGA [233], LEDITS [234], OIR-Diffusion [235] , leaf, text width=23em ] ] ]